OpenStudy (anonymous):

you are getting ready to make a 640-mile car trip . In general , your car can drive 160 miles on 5 gallons of gasoline. How many gallons of gasoline will you use for the trip? you started out with 4 gallons of gasoline in your car and gasoline is $2.29 per gallon. How much money will you spend on gasoline on the trip ?

OpenStudy (anonymous):

\[\frac{640*5}{160}=20 \](20-4) 2.29 = 36.64

OpenStudy (tylerd):